Italian Queue 2014 Prev Next

Mar 04-05

Observer: M. Fumana, R. Speziali, S. Bisogni
Telescope Operator: Geno
Support Astronomer: Dave (on call)
Instrument: MODS

Summary

At the beginning of the night we stay close due to high humidity. Sky covered. We could perform 4 hours of observation only. We had to stop due to high humidity.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Gilli Q2 J1048+46 red grating 1h airmass 1.14-
Gilli Q2 J1048+46 red grating 1h airmass ???
Vanzella C RXJ1532 dual grating 0.5h airmass 1.0
Standard BD+33d2642 dual grating - airmass 1.0, performed in deep twilight
Standard BD+33d2642 red grating - airmass 1.0, performed in deep twilight

Details

03:50 The TO opens the dome
  • Waiting for collimation
  • Mirror too warm (DT=4C)

  • GILLI

04:40 Slew to Gilli Q2
  • acqMODS 4Q2_J1048_M2_MOS.acq.
  • Seeing 1.8"
  • modsAlign mods.7.mms mods1r.20140305.0001.fits mods1r.20140305.0002.fits

05:01 execMODS 4Q2_J1048_M2_mos_1.obs mods1r 0004-0007
  • the average seeing is 1.7" at the beginning of the integration
  • the average seeing is nearly 2" during the first part of the integration of the second frame
  • active optics errors, we go on integrating with the last image.

06:14 the TO cancels the preset and tries to collimate, so we have to acquire again. After the collimation the seeing is really better (average seeing 1.2").

06:18 acqMODS 4Q2_J1048_M2_MOS.acq
  • modsAlign mods.7.mms mods1r.20140305.0008.fits mods1r.20140305.0009.fits

06:31 execMODS 4Q2_J1048_M2_mos_2.obs mods1r 0011-0014
  • the average seeing is ~1" at the beginning of the integration

07:40 TO CLOSES the dome due to humidity

11:45 TO opens the dome again and tries to collimate
  • humidity 40-45%
  • average seeing ~1.2"

  • VANZELLA C

11:50 acqMODS rxj1532_mos.acq
  • modsAlign rxj1532_mos.mms mods1r.0015 mods1r.0016
  • sent a -0.42" offset in y by hand

12:11 execMODS rxj1532_mos_1.obs mods1b 0001 mods1r 0019
  • average seeing 1.1" at the beginning of the integration

12:4 we stop the integration: twilight

  • STANDARD

12:46 acqMODS bd332642.acq
  • modsAlign -r mods1r.20140305.0020.fits

12:54 execMODS bd332642.obs
  • average seeing 1.4 at the beginning of the integration
  • dual grating mods1b 0002-0004 mods1r 0022-0024
  • lost the guide, we are tracking only
  • red only mods1r 0025-0027

01:22 The TO closes the dome

Topic revision: r8 - 06 Mar 2014, MarcoFumana
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback